OpenStudy (anonymous):

What is the distance from (−4, 0) to (2, 5)? Round your answer to the nearest hundredth.

imqwerty (imqwerty):

distance =((-4-2)^2 + (2-5)^2)^(1/2) = (36+9)^(1/2) =(45)^(1/2) =6.7082 rounding to nearest hundredth = 6.7